# The specification approach to Equilibrium States for Parabolic Rational Maps

### v1

@@ Abstract.
We develop the specification and orbit-decomposition approach to equilibrium states for parabolic rational maps of the Riemann Sphere. Our result extends the well-known results on uniqueness of equilibrium states in this setting, notably the results of Denker, Przytycki and Urbański. We extend the class of potentials from Hölder to those with the Bowen property on ‘good orbits’ and present an improvement on the pressure gap condition. Our pressure gap condition is sharp in the class of potentials we consider. In the family of geometric potentials, our approach gives a simple new proof of uniqueness of equilibrium states up to the phase transition that occurs at the Hausdorff dimension of the Julia set. For Hölder potentials, we show that our equilibrium state has positive entropy. This shows a posteriori that our pressure gap is equivalent to the a priori stronger condition that the potential is hyperbolic and this yields strong statistical properties for the equilibrium state.
@@ 1. Introduction
We develop the specification and orbit-decomposition approach to equilibrium states in the setting of parabolic rational maps. A rational map $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is an analytic endomorphism of the Riemann sphere, $\widehat{\mathbb{C}}$. For $z\in{\mathbb{C}}$, $f$ has the form $f(z)=P(z)/Q(z)$ where $P(z)$ and $Q(z)$ are nonzero relatively prime polynomials. For a rational map $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$, we denote its Julia set by $J(f)$. We consider parabolic rational maps $f$. These are the rational maps which have no critical points in the Julia set and have at least one rationally indifferent periodic point, i.e. a point $p$ such that $f^{n}(p)=p$ for some $n>0$ and $(f^{n})^{\prime}(p)$ is a root of unity. Let $\Omega$ denote the set of rationally indifferent periodic points for $f$. The set $\Omega$ is finite and contained in $J(f)$ and is in the closure of the forward orbit of a critical point. The system $(J(f),f)$ is a well-known example of a non-uniformly hyperbolic dynamical system.
@@
Thermodynamic formalism for rational maps was first developed for the measure of maximal entropy in papers by Freire, Lopes, Lyubich and Mañé [19, 23, 24]. Equilibrium states and related ergodic theory of this setting were developed in a series of papers by Denker, Przytycki, Urbański and co-authors, notably in [16, 14]. They proved in [14] that if $f$ is a rational map with $\deg(f)\geq 2$ and $\varphi:J(f)\rightarrow{\mathbb{R}}$ is a Hölder continuous potential satisfying the pressure gap condition $\sup_{J}\varphi:=\sup_{z\in J}\varphi(z)<P(\varphi)$, then there exists a unique equilibrium state. Here, $P(\varphi)$ is the topological pressure of the potential $\varphi$ on the compact dynamical system $(J(f),f)$. We say that $\varphi$ is a hyperbolic potential if there exists $n\geq 1$ such that
(1.1) $\sup_{J}\frac{1}{n}S_{n}\varphi:=\sup_{z\in J}\frac{1}{n}S_{n}\varphi(z)<P(\varphi).$
@@
The hyperbolicity condition is a priori weaker than the pressure gap condition of [14]. By [22, Proposition 3.1], if $\varphi$ satisfies (1.1) there is a cohomologous function $\tilde{\varphi}$ which satisfies $\sup_{J}\tilde{\varphi}<P(\varphi)$. It thus follows from [22] that a hyperbolic Hölder potential has a unique equilibrium state.
The papers [28, 1, 17, 20] further explored the ergodic theory of the parabolic case. A wide class of non-uniformly hyperbolic rational maps, mostly without parabolics, were studied by Przytycki and Rivera-Letelier [30]. There is an extensive literature on thermodynamic formalism in complex dynamics including [8, 25, 29, 30, 35, 38]. There has been a recent resurgence in interest in this area in papers such as [2, 3, 4, 5].
@@
which is the maximal average of $\varphi$ along any rationally indifferent periodic orbit. We consider continuous potentials $\varphi:J(f)\to{\mathbb{R}}$ which are ‘Bowen on good’, which means that $\varphi$ has the Bowen property on ${\mathcal{G}}(\eta)$ for every $\eta>0$. We define this property precisely in §2. Roughly, it is the property that there is a bound on variations of Birkhoff sums of $\varphi$ on those orbit segments ${\mathcal{G}}(\eta)$ that spend a positive proportion of time $\eta$ away from the parabolic cycles. We prove the following result.
@@
Let $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a parabolic rational map of degree $d\geq 2$ and $\Omega=\{w_{1},\dots,w_{k}\}$ be the set of its rationally indifferent periodic points. If $\varphi:J(f)\to{\mathbb{R}}$ has the Bowen property on ${\mathcal{G}}(\eta)$ for every $\eta>0$ and $A(\Omega,\varphi)<P(J(f),\varphi)$, then $\varphi$ has a unique equilibrium state w.r.t. the dynamical system $(J(f),f)$ and it is fully supported on $J(f)$.
We show that every Hölder continuous potential $\varphi$ has the Bowen property on ${\mathcal{G}}(\eta)$. The Hölder class is given by any smooth metric on $J(f)$. For parabolic rational maps, the Fatou set is always non-empty, so we can analytically conjugate $f$ to a map $g$ whose Julia set $J(g)$ is contained in $\mathbb{C}$. Then we can take our smooth metric to be the Euclidean metric on $J(g)$.
In our setting, the pressure gap $A(\Omega,\varphi)<P(\varphi)$ is a sharp criterion to have a unique equilibrium state which is fully supported. This is because each periodic orbit in $\Omega$ corresponds to an invariant measure, so if the pressure gap fails, then at least one of these measures is an equilibrium state. We emphasize that the ‘obstruction to uniqueness’ term $A(\Omega,\varphi)$ in our pressure gap is local, whereas the corresponding term in (1.1) is a global quantity.
@@
We emphasize that the corollary is not immediate from the hyperbolicity criterion (1.1) since it can happen that $\sup_{J}\frac{1}{n}S_{n}\varphi_{t}>0$ for $t>0$, so we cannot simply use the fact that $P(\varphi_{t})>0$ for all $t<h$. This demonstrates the utility of our improved criterion. We could not find a clear statement of the corollary in the literature, although we expect it is folklore. It follows for $t\in[0,h)$ as a corollary of Haydn and Isola’s results on analyticity of the pressure function [21]. We further discuss this application and previous results for geometric potentials in §4.
We now investigate further properties of the equilibrium state. Our key result is the following.
@@
Suppose we are in the setting of Theorem 1.1. The unique equilibrium state has positive entropy.
For a Hölder potential, this result has strong consequences since it implies that the potential is hyperbolic. Via the results of [22], we can then apply the results of Szostakiewicz, Urbanski and Zdunik [34] to obtain the following.
@@
Let $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a parabolic rational map of degree $d\geq 2$ and $\Omega=\{w_{1},\dots,w_{k}\}$ be the set of its rationally indifferent periodic points. Let $\varphi:J(f)\rightarrow{\mathbb{R}}$ be a Hölder potential. The following are equivalent:
@@
In particular, if $A(\Omega,\varphi)<P(\varphi)$, then the unique equilibrium state satisfies the strong statistical properties provided by [34]: exponential decay of correlations, central limit theorem and law of iterated logarithms.
@@
Our proof of Theorem 1.1 is based on the specification and orbit-decomposition approach to thermodynamic formalism, and is obtained by verifying criteria developed by Climenhaga and the second-named author. Theorem B follows from work of Call and the second-named author and follows from lifting the orbit decomposition to the product system $(J(f)\times J(f),f\times f)$. To the best of our knowledge, this is the first time that a specification-based approach to equilibrium states has been developed in a complex dynamics context [^1] We note that specification ideas were used in [31, Appendix A] to connect the topological Collet-Eckman condition and uniform hyperbolicity on periodic orbits. It was observed there that these ideas have independent interest.. Our proof thus has methodological interest, and it is intended to lay the foundations for further exploration of the specification and orbit-decomposition techniques for rational maps and beyond. Existing results are usually based on the powerful and well-developed machinery of transfer operators and inducing schemes, which have had great success in this area. We anticipate that further development of the specification approach will provide a complementary toolkit which has some advantages.
@@ 2.1. Thermodynamic Formalism
The above result was proved by Climenhaga and the second named author in [13, Theorem 5.5] for homeomorphisms with a non-uniform expansivity assumption. The non-invertible case with a non-uniform positive expansivity assumption is essentially the same proof, with some simpler book-keeping. In the measure of maximal entropy case, the non-invertible case was treated with full details in [12].
@@ 2.4. Metric expansion properties and the Milnor metric
We now discuss metric expansion properties, and we refer the reader to [26, Appendix B] for more details. First, we discuss some natural metrics. For a parabolic rational map, the hyperbolic metric is only well-defined and expanding on $f^{-1}(\widehat{\mathbb{C}}\setminus P(f))$. We thus cannot use the hyperbolic metric for an expanding metric at the parabolic points or their preimages. We could consider the Euclidean metric on $J(f)$. Since $J(f)$ contains no critical points, we know that there exists $c>0$ so that $|f^{\prime}(z)|\geq c>0$ for all $z\in J$, but we do not have any better global lower bound. However, as a consequence of the Leau-Fatou Flower Theorem, in a punctured neighborhood of a parabolic fixed point the Euclidean derivative is strictly greater than one. Milnor proved that a parabolic rational map $f$ admits a metric such that $f$ is distance-increasing on $J(f)$. The idea is to combine the desirable properties of the hyperbolic and Euclidean metrics. Milnor’s metric is defined piecewise, as a rescaling of the Euclidean metric near the rationally indifferent periodic points and the hyperbolic metric away from them. This construction is the key point which proves the following lemma.
@@
To simplify our estimates, we proceed under the assumption that $\Omega(f)$ has only fixed points. The proof that $d_{\alpha}$ is distance-increasing on $J(f)$ is in [26, Lemma B.3]. We adapt the proof to show that $f$ is uniformly distance-expanding (w.r.t $d_{\alpha}$) on $K(\alpha)$ follows part of Milnor’s proof in [26, Lemma B.3].
@@ 4. Family of geometric potentials
Since there are no critical points in $J(f)$, the functions $\varphi_{t}:J(f)\to{\mathbb{R}}$ are well-defined and Hölder continuous. Note that the definition of $\varphi$ depends on the choice of metric. If we define the geometric potential using a different metric, it follows from (2.2) that we differ by a coboundary, so the pressure and the set of equilibrium states are the same. Let $h=\text{HD}(J(f))$ denote the Hausdorff dimension of $J(f)$. Our claim is that $\varphi_{t}$ has a unique equilbrium state for $t<h$.
@@
As discussed in the introduction, it is not clear how to directly check the hyperbolicity criterion since we do not have a sign on $\sup_{J}\frac{1}{n}S_{n}\varphi_{t}$ in the Euclidean metric. This can be seen if we arrange to have a point $z\in f^{-1}\Omega$ with $|f(z)|<1$ and $f^{2}z=f(z)$. We see that $S_{n}\varphi_{t}(z)>0$ for all $n$.
@@
A classical approach to this corollary would be to use the existence of a Markov partition [32, 1] for a parabolic rational map and recode to a countable-state topological Markov chain with good distortion properties. This approach was taken when the analogue of Corollary 1.2 was proved for the Manneville-Pomeau map by Sarig in [33, Claim 2]. We would not be surprised if this approach has been developed for parabolic rational maps previously, but we could not find it in the literature; the paper [1] develops related tools but does not investigate equilibrium states. We also mention the results of Przytycki and Rivera-Letelier [30] who used inducing schemes to study the potentials $t\log|f^{\prime}(z)|$ for an interval of values around 0 for a wide class of non-uniformly hyperbolic rational maps allowing critical points in the Julia set, although their paper excludes parabolics except for quadratic map with real coefficients.
By results of Walters [37], uniqueness of equilibrium states for the interval $(-\infty,h)$ holds if and only if the pressure function $t\to P(\varphi_{t})$ is $C^{1}$ on this interval. There is a phase transition at $t=h$ where uniqueness of the equilibrium state and smoothness of the pressure function fails. Our argument thus gives a simple justification for the famous picture of a non-uniformly hyperbolic pressure function in this setting. Haydn and Isola [21] showed by hard analysis of the spectrum of the transfer operator that the pressure function is analytic for $t\in[0,h)$ which thus provides an earlier indirect proof of our corollary.

### v2

@@ Abstract.
We develop the specification and orbit-decomposition approach to equilibrium states for parabolic rational maps of the Riemann Sphere. Our result extends the well-known results on uniqueness of equilibrium states in this setting, notably the results of Denker, Przytycki and Urbański. We extend the class of potentials from Hölder to those with the Bowen property on ‘good orbits’ . We obtain uniqueness of the equilibrium state for potentials satisfying a pressure gap condition which is sharp in the class of potentials we consider. We show that our equilibrium state has the $K$-property, and in particular it has positive entropy. When the potential is Hölder, the theory of equilibrium states is already highly developed. Nevertheless, several interesting results on equilibrium states for Hölder potentials follow readily from our approach. In the family of geometric potentials, we obtain a simple proof of uniqueness of equilibrium states up to the phase transition that occurs at the Hausdorff dimension of the Julia set. For Hölder potentials on parabolic rational maps, we show that hyperbolicity of the potential is equivalent to having a unique equilibrium state which is fully supported. This does not appear to have been stated in the literature before, although it may be considered folklore.
@@ 1. Introduction
We develop the specification and orbit-decomposition approach to equilibrium states in the setting of parabolic rational maps. Recent literature on the specification approach is split into papers which build general theory for equilibrium states [17, 13], and papers which apply the general theory to obtain results for specific classes of dynamical systems using arguments and structures specialized to these systems [11, 14, 35]. This paper falls into the second category and is focused on obtaining refined uniqueness of equilbrium states results for parabolic rational maps. A rational map $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is an analytic endomorphism of the Riemann sphere, $\widehat{\mathbb{C}}$. For $z\in{\mathbb{C}}$, $f$ has the form $f(z)=P(z)/Q(z)$ where $P(z)$ and $Q(z)$ are nonzero relatively prime polynomials. For a rational map $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$, we denote its Julia set by $J(f)$. We consider parabolic rational maps $f$. These are the rational maps which have no critical points in the Julia set and have at least one rationally indifferent periodic point, i.e. a point $p$ such that $f^{n}(p)=p$ for some $n>0$ and $(f^{n})^{\prime}(p)$ is a root of unity. Let $\Omega$ denote the set of rationally indifferent periodic points for $f$. The set $\Omega$ is finite and contained in $J(f)$ and is in the closure of the forward orbit of a critical point. The system $(J(f),f)$ is a well-known example of a non-uniformly hyperbolic dynamical system.
@@
Thermodynamic formalism for rational maps was first developed for the measure of maximal entropy in papers by Freire, Lopes, Lyubich and Mañé [25, 30, 31]. Equilibrium states and related ergodic theory of this setting were developed in a series of papers by Denker, Przytycki, Urbański and co-authors, notably in [22, 20] with an important improvement contributed by Inoquio-Renteria and Rivera-Letelier in [28]. The landmark result is that if $f$ is a rational map with $\deg(f)\geq 2$ and $\varphi:J(f)\rightarrow{\mathbb{R}}$ is a hyperbolic Hölder continuous potential, then there exists a unique equilibrium state. Here, we say that $\varphi$ is a hyperbolic potential if there exists $n\geq 1$ such that
(1.1) $\sup_{J}\frac{1}{n}S_{n}\varphi:=\sup_{z\in J}\frac{1}{n}S_{n}\varphi(z)<P(\varphi),$
@@
where $P(\varphi)$ is the topological pressure of the potential $\varphi$ on the compact dynamical system $(J(f),f)$. In the seminal result of [20], uniqueness of equilibrium states was proved using the a priori stronger hypothesis that $\sup_{J}\varphi:=\sup_{z\in J}\varphi(z)<P(\varphi)$. By [28, Proposition 3.1], if $\varphi$ satisfies (1.1) there is a cohomologous function $\tilde{\varphi}$ which satisfies $\sup_{J}\tilde{\varphi}<P(\varphi)$. It thus follows from [28] that a hyperbolic Hölder potential has a unique equilibrium state.
Recall that a potential is hyperbolic if and only if all equilibrium states for the potential have positive entropy. The failure of uniqueness is often associated with the presence of a zero entropy equilibrium state at a phase transition. While hyperbolicity is a mild and natural condition, it is not perfect from the point of view of the uniqueness theory since it a priori rules out the likely source of non-uniqueness which is the possibility of zero entropy equilibrium states. The room for growth in this theory is to find natural criteria for a potential to have a unique equilibrium state without assuming that the potential is hyperbolic or to establish weaker criterion to verify hyperbolicity.
@@
which is the maximal average of $\varphi$ along any rationally indifferent periodic orbit. Our focus is on Hölder continuous potentials, with the Hölder class given by any smooth metric on $J(f)$. We prove the following result.
@@
Let $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a parabolic rational map of degree $d\geq 2$ and $\Omega=\{w_{1},\dots,w_{k}\}$ be the set of its rationally indifferent periodic points. If $\varphi:J(f)\to{\mathbb{R}}$ is Hölder and $A(\Omega,\varphi)<P(J(f),\varphi)$, then $\varphi$ has a unique equilibrium state w.r.t. the dynamical system $(J(f),f)$ and it is fully supported on $J(f)$.
Theorem 1.1 holds when the Hölder assumption on $\varphi$ is weakened to the technical dynamical regularity condition that $\varphi$ is ‘Bowen on good orbit segments’. We define this property in §2 and we verify that the ‘Bowen on good’ property holds for Hölder potentials in §3.2. We state the more general version of our result as Theorem 3.11.
In our setting, the pressure gap $A(\Omega,\varphi)<P(\varphi)$ is a sharp criterion to have a unique equilibrium state which is fully supported. This is because each periodic orbit in $\Omega$ corresponds to an invariant measure, so if the pressure gap fails, then at least one of these measures is an equilibrium state. We emphasize that the ‘obstruction to uniqueness’ term $A(\Omega,\varphi)$ in our pressure gap is local, whereas the corresponding term in the hyperbolicity condition (1.1) is a global quantity.
@@
We could not find an explicit statement of the corollary in the literature, but we emphasize that it can be obtained readily from a criteria of Inoquio-Renteria and Rivera-Letelier which we discuss in §4. We now investigate further properties of the equilibrium state. Our key result is the following.
@@
Assume the setting of Theorem 1.1. The unique equilibrium state has the $K$-property. In particular, the equilibrium state has positive entropy.
For a Hölder potential, this result has strong consequences since positive entropy implies that the potential is hyperbolic. Via the results of [28], we can then apply the results of Szostakiewicz, Urbanski and Zdunik [44] to obtain the following.
@@
Let $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a parabolic rational map of degree $d\geq 2$ and $\Omega=\{w_{1},\dots,w_{k}\}$ be the set of its rationally indifferent periodic points. Let $\varphi:J(f)\rightarrow{\mathbb{R}}$ have the Bowen property on ${\mathcal{G}}(\eta)$ for every $\eta>0$. The following are equivalent:
@@
In particular, if $\varphi$ is Hölder and $A(\Omega,\varphi)<P(\varphi)$, then the unique equilibrium state satisfies the strong statistical properties provided by [44]: exponential decay of correlations, central limit theorem and law of iterated logarithms.
From (2), if $\varphi$ is not hyperbolic, we must have $A(\Omega,\varphi)=P(\varphi)$ and thus there is an equilibrium state which is a periodic orbit measure in $\Omega$. Thus, the hyperbolicity property is not only sufficient for, but equivalent to, the property that there is a unique equilibrium state and the equilibrium state is fully supported. Although this result may be folklore, we emphasize that it follows readily from our approach and has not previously appeared in the literature [^1] After completing this work, Juan Rivera-Letelier shared with us an elegant unpublished argument to show that $A(\Omega,\varphi)<P(\varphi)$ implies hyperbolicity of the potential based on direct analysis of Lyapunov exponents and facts from complex analysis. Analogous results have been proved for maps of the interval using different techniques, see e.g. [19, Appendix A], [10, Corollary 3.8]..
Corollary 1.5.
Let $f:\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a parabolic rational map of degree $d\geq 2$. Let $\varphi:J(f)\rightarrow{\mathbb{R}}$ be Hölder continuous. Then there is a unique equilibrium state $\mu$, with $\mu$ fully supported on $J(f)$, if and only if $f$ is hyperbolic.
Our proof of Theorem 1.1 is based on the specification and orbit-decomposition approach to thermodynamic formalism, and is obtained by verifying criteria developed by Climenhaga and the second-named author. These techniques have previously been applied mostly in invertible or continuous-time settings, but the general machinery applies equally to endomorphisms, as established in the original papers in this direction, e.g. [17]. Theorem B follows from work of Call and the second-named author and follows from lifting the orbit decomposition to the product system $(J(f)\times J(f),f\times f)$.
@@
To the best of our knowledge, this is the first time that a specification-based approach to equilibrium states has been developed in a complex dynamics context [^2] We note that specification ideas were used in [40, Appendix A] to connect the topological Collet-Eckman condition and uniform hyperbolicity on periodic orbits. It was observed there that these ideas have independent interest.. Our proof thus has methodological interest, and it is intended to lay the foundations for further exploration of specification techniques in complex dynamics. Existing results are usually based on the powerful and well-developed machinery of transfer operators and inducing schemes, which have had great success in this area. We anticipate that further development of the specification approach in complex dynamics will provide a complementary toolkit which has some advantages.
We briefly discuss further literature surrounding this area. There has been a recent resurgence in interest in thermodyamic formalism and complex dynamics in papers such as [3, 4, 5, 6]. In addition to the key references for general rational maps already discussed, particularly [20, 44], the papers [36, 1, 23, 26] extensively develop the ergodic theory of the parabolic case, focusing on the natural measure and its dimension theory. A wide class of non-uniformly hyperbolic rational maps, mostly without parabolics, were studied by Przytycki and Rivera-Letelier [39]. There is an extensive literature on thermodynamic formalism in complex dynamics including [9, 32, 38, 39, 46, 49]. We particularly recommend the textbook [45] for a thorough treatment of advanced topics in this area.
There has been extensive progress in recent years in the general theory of equilibrium states. In addition to the framework provided by specification techniques, there has been remarkable progress in smooth settings based on countable-state symbolic dynamics. The seminal work of Buzzi, Crovisier and Sarig [12] implies that a hyperbolic potential for a smooth diffeomorphism of a surface has a unique equilibrium state. This theory has been developed for general smooth endomorphisms in [2, 29]. The coding is constructed to capture typical points for all measures with exponent above some fixed threshold, and this can be used to show that there is at most one hyperbolic equilibrium state. To use such codings for uniqueness results among all equilibrium states, the possibility of zero exponent equilibrium states must be excluded a priori. Thus, the symbolic techniques do not yield new information for Hölder potentials for rational maps.
@@ 2.1. Thermodynamic Formalism
The above result was proved by Climenhaga and the second named author in [18, Theorem 5.5] for homeomorphisms with a non-uniform expansivity assumption. The non-invertible case with a non-uniform positive expansivity assumption is essentially the same proof, with some simpler book-keeping. In the measure of maximal entropy case, the non-invertible case was treated with full details in [17].
@@ 2.4. Metric expansion properties and the Milnor metric
We now discuss metric expansion properties, and we refer the reader to [33, Appendix B] for more details. First, we discuss some natural metrics. For parabolic rational maps, the Fatou set is always non-empty, so we can analytically conjugate $f$ to a map $g$ whose Julia set $J(g)$ is contained in $\mathbb{C}$. Then the Euclidean metric is a smooth metric on $J(g)$. For a parabolic rational map, the hyperbolic metric is only well-defined and expanding on $f^{-1}(\widehat{\mathbb{C}}\setminus P(f))$. We thus cannot use the hyperbolic metric for an expanding metric at the parabolic points or their preimages. We could consider the Euclidean metric on $J(f)$. Since $J(f)$ contains no critical points, we know that there exists $c>0$ so that $|f^{\prime}(z)|\geq c>0$ for all $z\in J$, but we do not have any better global lower bound. However, as a consequence of the Leau-Fatou Flower Theorem, in a punctured neighborhood of a parabolic fixed point the Euclidean derivative is strictly greater than one. Milnor proved that a parabolic rational map $f$ admits a metric such that $f$ is distance-increasing on $J(f)$. The idea is to combine the desirable properties of the hyperbolic and Euclidean metrics. Milnor’s metric is defined piecewise, as a rescaling of the Euclidean metric near the rationally indifferent periodic points and the hyperbolic metric away from them. This construction is the key point which proves the following lemma.
@@
To simplify our estimates, we proceed under the assumption that $\Omega(f)$ has only fixed points. The proof that $d_{\alpha}$ is distance-increasing on $J(f)$ is in [33, Lemma B.3]. We adapt the proof to show that $f$ is uniformly distance-expanding (w.r.t $d_{\alpha}$) on $K(\alpha)$ follows part of Milnor’s proof in [33, Lemma B.3].
@@ 4. Family of geometric potentials
Since there are no critical points in $J(f)$, the functions $\varphi_{t}:J(f)\to{\mathbb{R}}$ are well-defined and Hölder continuous. Note that the definition of $\varphi$ depends on the choice of metric. If we define the geometric potential using a different metric, it follows from (2.2) that we differ by a coboundary, so the pressure and the set of equilibrium states are the same. Let $h=\text{HD}(J(f))$ denote the Hausdorff dimension of $J(f)$. Our claim is that $\varphi_{t}$ has a unique equilbrium state for $t<h$. This is not a new result, and we discuss other ways to obtain it below, but it is instructive that it follows easily from our criterion.
@@
The proof can be obtained another way using existing results by checking that $\varphi_{t}$ is hyperbolic for $t<h$. Note that hyperbolicity is not immediate since we do not have a sign on $\sup_{J}\frac{1}{n}S_{n}\varphi_{t}$ in the Euclidean metric. For example, there are parabolic rational maps for which there exists $z\in f^{-1}\Omega$ with $|f^{\prime}(z)|<1$ and $f(z)$ a fixed point. We see that $S_{n}\varphi_{t}(z)>0$ for all $n$. Thus we cannot simply use the fact that $P(\varphi_{t})>0$ for all $t<h$ to conclude that $\varphi_{t}$ is hyperbolic for $t<h$. Nevertheless, hyperbolicity can easily be verified from the existing literature. Przytycki showed in [37] that any invariant measure has non-negative Lyapunov exponent, see [41, Appendix A] for a short recent proof. Thus for $t>0$ and any measure $\mu$, we have $t\int \log|f’|d\mu\leq 0$. A potential is hyperbolic if and only if $\sup\int\varphi d\mu<P(\varphi)$, see [28, Proposition 3.1]. If $t<h$, we thus have $\sup\int\varphi_{t}d\mu\leq 0<P(\varphi_{t})$ and it follows that $\varphi_{t}$ is hyperbolic.
Remark.
Another approach would be to use the existence of a Markov partition [42, 1] for a parabolic rational map and recode to a countable-state topological Markov chain with good distortion properties. This approach was taken when the analogue of Corollary 1.2 was proved for the Manneville-Pomeau map by Sarig in [43, Claim 2]. We would not be surprised if this approach has been developed for parabolic rational maps previously, but we could not find it in the literature; the paper [1] develops related tools but does not investigate equilibrium states. We also mention that Przytycki and Rivera-Letelier [39] used inducing schemes to study the potentials $t\log|f^{\prime}(z)|$ for an interval of values around 0 for a wide class of non-uniformly hyperbolic rational maps allowing critical points in the Julia set, although they exclude parabolics except for quadratic map with real coefficients.
@@
By results of Walters [48], uniqueness of equilibrium states for $(-\infty,h)$ holds if and only if the pressure function $t\to P(\varphi_{t})$ is $C^{1}$ on this interval. There is a phase transition at $t=h$ where uniqueness of the equilibrium state and smoothness of the pressure function fails. Our argument thus gives a simple justification for the famous picture of a non-uniformly hyperbolic pressure function in this setting. We remark that Haydn and Isola [27] showed by hard analysis of the spectrum of the transfer operator that the pressure function is analytic for $t\in[0,h)$ which thus provides the earliest indirect proof that we know of for our corollary.